What is the slope of the line represented by the equation below?
2y = x +3
1
1/2
3/2
3

Answer by stanbon(75887)

What is the slope of the line represented by the equation below?
2y = x +3
----
Write in slope-intercept form:::
y = (1/2)x + (3/2)
----
slope = 1/2
=============
Cheers,
Stan H.
=============
1
1/2
3/2
3